2. Determine the values of E°cell and ΔG° for the following reactions.

(a) O2(g) + 4 I–(aq) + 4 H+(aq) → 2 H2O(l) + 2 I2(s) E°cell _________ ΔG°__________

(b) 4 IO3–(aq) + 4 H+(aq) → 2 I2(s) + 2 H2O(l) + 5 O2(g) E°cell _________ ΔG°_________
Answer #1

a) Oxidation half reaction : 4 I - (aq) \rightarrow 2 I 2 (s) + 4 e -

Reduction half equation : O 2 (g) + 4 H + (aq) + 4 e -\rightarrow 2 H2O (l)  

Overall reaction : O 2 (g) + 4 H + (aq) + 4 I - (aq) \rightarrow   2 H2O (l) +  2 I 2 (s)

E 0 cell for the overall reaction is E 0 cell = E 0 reduction - E 0 oxidation  

\therefore E 0 cell = 1.229 V - 0.535 V = 0.694 V

We have relation , \Delta G 0 = - n F  E 0 cell

Where n = no. of moles of electrons transferred in the overall reaction.

We have, n = 4 , F = 96485 C / mol &  E 0 cell = 0.694 V

\therefore\DeltaG 0 = - 4 \times 96485 \times 0.694 =- 267842 J / mol =- 267.8 k J / mol

ANSWER :  E 0 cell = 0.694 V & \Delta G 0 = - 267.8 k J / mol

b)

Oxidation half reaction : 10 H2O (l)    \rightarrow 5 O 2 (g) + 20 H + (aq) + 20 e -

Reduction half equation : 4 IO3- (aq) + 24 H + (aq) + 20 e -\rightarrow 2 I 2 (s) + 12 H2O (l)

Overall reaction : 4 IO3- (aq) + 4 H + (aq) \rightarrow 5 O 2 (g) +  2 I 2 (s) + 2 H2O (l)  

E 0 cell for the overall reaction is E 0 cell = E 0 reduction - E 0 oxidation  

\therefore E 0 cell = 1.210 V - 1.229 V = - 0.019 V

We have relation , \Delta G 0 = - n F  E 0 cell

Where n = no. of moles of electrons transferred in the overall reaction.

We have, n =20 , F = 96485 C / mol &  E 0 cell = - 0.019 V

\therefore\DeltaG 0 = - 20 \times 96485 \times (- 0.019) = 36664 J / mol =36.664 k J / mol

ANSWER :  E 0 cell = - 0.019 V & \Delta G 0 = 36.664 k J / mol